# From attention to profit: quantitative trading strategy based on transformer

### v1

@@ Appendix A Detailed backtest results between 100 factors and the transformer factor
  • Max Drawdown (MD) means the potential worst-case scenario or the most extreme possible loss.
  • Portfolio volatility ($\sigma_{p}$) is used to measure the risk of a strategy, where is positively correlated to the risk of a strategy.
Factor AR (%) AER (%) WR (%) SR Alpha Beta MD (%) $\sigma_{p}$
ARBR -1.47 0.32 54.7 -0.251 -0.002 0.914 34.46 0.221
AR -5.87 -4.37 53.7 -0.406 -0.048 0.981 47.84 0.260
ATR14 -2.12 -0.37 62.7 -0.250 0.006 1.160 46.55 0.250
ATR6 -3.94 -2.31 61.2 -0.328 -0.014 1.175 50.11 0.253
BBIC 1.17 3.14 50.9 -0.109 0.033 1.038 32.64 0.258
BR -6.91 -5.47 53.6 -0.467 -0.064 0.934 51.72 0.254
CCI10 -18.65 -17.99 43.3 -1.402 -0.261 0.812 72.04 0.220
CCI15 -17.70 -16.98 44.2 -1.293 -0.240 0.832 71.43 0.223
CCI20 -15.52 -14.66 46.3 -1.051 -0.196 0.868 69.12 0.235
CR20 -4.27 -2.67 54.8 -0.329 -0.032 0.927 49.92 0.263
DAVOL10 -7.44 -6.04 54.1 -0.504 -0.071 0.920 45.45 0.248
DAVOL20 -10.95 -9.78 54.4 -0.702 -0.119 0.919 48.96 0.246
DAVOL5 -9.54 -8.28 51.8 -0.612 -0.097 0.950 46.41 0.250
EMA5 2.77 4.85 48.3 -0.047 0.049 1.041 30.70 0.259
EMAC10 0.27 2.17 50.6 -0.144 0.024 1.037 31.42 0.258
EMAC12 1.58 3.57 50.6 -0.093 0.037 1.032 31.76 0.258
EMAC20 6.01 8.29 54.1 0.069 0.079 1.044 32.43 0.260
EMAC26 3.83 5.97 54.1 -0.008 0.059 1.046 34.87 0.261
EMAC120 1.74 3.75 48.3 -0.088 0.034 0.960 35.81 0.254
Kurtosis20 -2.96 -1.27 52.5 -0.395 -0.027 0.772 27.71 0.182
Kurtosis60 -0.28 1.59 53.9 -0.240 -0.001 0.716 22.07 0.179
Kurtosis120 -4.12 -2.51 49.0 -0.468 -0.043 0.708 28.28 0.182
MAC5 0.54 2.46 46.9 -0.134 0.026 1.035 30.03 0.257
MAC10 -1.20 0.61 49.6 -0.206 0.008 1.039 34.15 0.256
MAC20 2.47 4.52 53.5 -0.059 0.046 1.043 34.58 0.257
MAC60 11.10 13.72 53.8 0.243 0.120 0.987 28.61 0.254
MAC120 3.11 5.21 47.7 -0.035 0.049 0.987 35.40 0.259
MACDC -11.12 -9.97 52.4 -0.674 -0.117 1.003 46.48 0.261
MASS -10.55 -9.35 51.0 -0.673 -0.115 0.886 48.34 0.249
MAWVAD 7.29 9.67 55.9 0.119 0.082 0.891 36.54 0.248
MFI14 -6.21 -4.73 51.5 -0.456 -0.056 0.918 45.65 0.241
PLRC6 -14.83 -13.92 47.8 -0.905 -0.178 0.960 66.24 0.259
PLRC12 -7.28 -5.87 53.3 -0.452 -0.064 1.012 52.30 0.273
PLRC24 -5.77 -4.26 58.1 -0.361 -0.044 1.034 53.28 0.289
PSY -9.11 -7.83 50.6 -0.660 -0.096 0.880 49.79 0.223
Price1M -13.55 -12.55 53.0 -0.760 -0.156 0.977 64.85 0.280
Price3M -6.69 -5.25 58.4 -0.380 -0.056 1.023 59.14 0.305
Price1Y 1.55 3.54 66.8 -0.075 0.041 1.116 60.25 0.326
ROC6 -16.13 -15.30 47.4 -0.982 -0.200 0.978 68.14 0.262
ROC20 -11.63 -10.51 54.7 -0.628 -0.123 1.014 60.28 0.291
ROC60 -6.86 -5.43 59.6 -0.382 -0.056 1.062 62.16 0.309
ROC120 2.36 4.41 66.9 -0.051 0.049 1.117 56.36 0.319
Skewness20 -5.91 -4.41 50.8 -0.538 -0.058 0.820 40.65 0.197
Skewness60 -6.47 -5.01 51.9 -0.595 -0.068 0.765 38.02 0.190
Skewness120 -4.37 -2.77 51.1 -0.459 0.044 0.743 36.01 0.192
TRIX5 -10.72 -9.54 52.5 -0.607 -0.110 1.009 56.07 0.280
TRIX10 -3.51 -1.85 58.4 -0.268 -0.019 1.008 52.92 0.292
TVMA20 -1.92 -0.16 64.9 -0.209 0.014 1.268 44.66 0.288
TVMA6 -0.75 1.09 62.4 -0.166 0.026 1.259 39.46 0.288
TVSTD20 -1.83 -0.06 58.5 -0.214 0.012 1.218 35.89 0.277
TVSTD6 -3.77 -2.13 53.4 -0.297 -0.010 1.202 46.57 0.272
VDEA -3.05 -1.37 57.6 -0.305 -0.023 0.857 37.01 0.239
VDIFF -7.93 -6.57 0.2 -0.533 -0.079 0.895 39.79 0.247
VEMA5 -0.41 1.46 57.9 -0.186 0.006 0.850 37.31 0.237
VEMA10 0.95 2.91 58.9 -0.128 0.019 0.844 35.51 0.236
VEMA12 1.01 2.96 59.3 -0.126 0.020 0.840 35.17 0.235
VEMA26 2.41 4.46 59.3 -0.068 0.033 0.838 31.03 0.231
VMACD -1.03 0.79 50.5 -0.216 0.001 0.875 33.96 0.235
VOL5 -8.12 -6.76 55.9 -0.452 -0.066 1.155 53.86 0.296
VOL10 -7.07 -5.64 58.3 -0.404 -0.053 1.145 50.65 0.298
VOL20 -5.43 -3.89 58.6 -0.337 -0.032 1.156 50.19 0.298
VOL60 0.42 2.34 58.1 -0.125 0.032 1.145 38.35 0.286
VOL120 -1.88 -0.12 53.0 -0.213 0.007 1.136 42.09 0.281
VOL240 -5.51 -3.98 47.0 -0.382 -0.038 1.075 43.53 0.265
VOSC -6.35 -4.88 51.2 -0.453 -0.056 0.944 40.40 0.246
VROC6 -16.30 -15.49 42.7 -1.179 -0.208 0.897 69.03 0.221
VROC12 -11.47 -10.34 46.7 -0.793 -0.127 0.915 51.62 0.228
VR -8.43 -7.10 52.8 -0.566 -0.083 0.932 48.57 0.244
VSTD10 -0.77 1.07 56.3 -0.205 0.001 0.843 37.75 0.234
VSTD20 -1.60 0.19 53.7 -0.245 -0.007 0.844 37.31 0.232
Variance20 -1.94 -0.18 59.9 -0.201 0.005 1.124 46.38 0.302
Variance60 9.96 12.51 65.2 0.168 0.122 1.178 30.63 0.313
Variance120 5.40 7.65 60.8 0.040 0.083 1.193 39.54 0.312
Volume1M -15.54 -14.68 20.7 -0.867 -0.189 0.991 63.68 0.285
WVAD -0.71 1.13 52.3 -0.199 0.002 0.836 42.38 0.238
arron_down_25
-7.35 -5.94 48.3 -0.593 -0.072 0.880 42.63 0.209
arron_up_25 -15.51 -14.65 46.4 -1.049 -0.195 0.870 66.73 0.235
bear_power -13.31 -12.30 50.8 -0.800 -0.154 0.934 59.58 0.261
beta 10.30 12.87 62.4 0.168 0.134 1.344 35.96 0.328
boll-up 2.19 4.23 54.5 -0.068 0.044 1.062 33.09 0.265
book_to_price 1.87 3.86 58.5 -0.117 0.031 0.889 26.73 0.181
bull_power -14.81 -13.91 51.1 -0.844 -0.176 0.984 65.99 0.278
earnings_yield -2.41 -0.68 52.6 -0.348 -0.029 0.627 25.87 0.188
growth 2.90 4.99 59.9 -0.045 0.052 1.068 37.35 0.244
leverage 2.05 4.07 53.1 -0.089 0.023 0.719 22.57 0.217
liquidity -2.71 -1.00 56.7 -0.243 -0.002 1.137 41.28 0.284
momentum 10.93 13.54 65.8 0.195 0.129 1.159 43.19 0.310
money_flow_20 -1.72 0.05 65.5 -0.202 0.016 1.269 44.10 0.288
price_no_fq 6.83 9.17 76.1 0.095 0.092 1.132 44.86 0.267
pull-down -2.28 -0.54 47.6 -0.317 -0.015 0.831 34.46 0.203
residual_volatility
4.13 6.30 62.3 0.003 0.062 1.048 41.34 0.269
sharpe_ratio_20
-11.88 -10.78 54.0 -0.663 -0.129 0.979 57.66 0.282
sharpe_ratio_60
-7.42 -6.02 58.8 -0.417 -0.065 1.029 65.19 0.300
sharpe_ratio_120
-0.32 1.55 65.7 -0.143 0.019 1.062 60.52 0.303
size -6.36 -4.89 51.8 -0.571 -0.058 0.906 36.00 0.196
turnover_volatility
-10.80 -9.62 51.8 -0.624 -0.108 1.075 57.05 0.274
1day_VPT -16.45 -15.65 44.8 -1.062 -0.211 0.904 70.43 0.248
1day_VPT_6
-6.36 -4.89 53.0 -0.416 -0.053 0.995 47.06 0.269
1day_VPT_12
-0.07 1.82 57.9 -0.146 0.020 1.033 46.51 0.278
25week_close
-11.51 -10.39 51.0 -0.859 -0.133 0.817 47.79 0.211
Trans_Month_1
17.35 19.43 57.8 0.915 0.162 0.588 18.35 0.161
Average -3.78 -2.15 53.71 -0.36 -0.03 0.98 44.88 0.26
CSI300 1.77 \ \ \ \ \ \ 0.197

### v2

@@ Appendix A Detailed backtest results between 100 factors and the quantformer factor
  • Max Drawdown (MD) means the potential worst-case scenario or the most extreme possible loss from the previous peak. For the trough value of a portfolio $V_{trough}$ and the peak value of the portfolio $V_{peak}$, the MDD is calculated in the following way:
    $MDD=\frac{V_{peak}-V_{trough}}{V_{peak}}$
  • Portfolio volatility ($\sigma_{p}$) is the standard deviation of the portfolio returns. $\sigma_{p}$ avoids the problem of negative deviations canceling with positive deviations and also penalizes larger deviations from the mean. It provides a kind of weighted average deviation in which large deviations carry more weight [58].
  • Sortino Ratio (STN) differentiates harmful volatility from total overall volatility by using the standard deviation of negative portfolio returns. This metric measures the performance of the investment relative to the downward risk.
[…]
Factor AR (%) AER (%) WR (%) SR Alpha Beta MD (%) $\sigma_{p}$ VaR (%) STN
ARBR -1.47 0.32 54.7 -0.251 -0.002 0.914 34.46 0.221 3.74 -0.375
AR -5.87 -4.37 53.7 -0.406 -0.048 0.981 47.84 0.260 3.94 0.612
ATR14 -2.12 -0.37 62.7 -0.250 0.006 1.160 46.55 0.250 4.51 -0.35
ATR6 -3.94 -2.31 61.2 -0.328 -0.014 1.175 50.11 0.253 4.48 -0.455
BBIC 1.17 3.14 50.9 -0.109 0.033 1.038 32.64 0.258 4.16 -0.155
BR -6.91 -5.47 53.6 -0.467 -0.064 0.934 51.72 0.254 3.80 -0.718
CCI10 -18.65 -17.99 43.3 -1.402 -0.261 0.812 72.04 0.220 2.79 -2.026
CCI15 -17.70 -16.98 44.2 -1.293 -0.240 0.832 71.43 0.223 2.90 -1.866
CCI20 -15.52 -14.66 46.3 -1.051 -0.196 0.868 69.12 0.235 3.28 -1.519
CR20 -4.27 -2.67 54.8 -0.329 -0.032 0.927 49.92 0.263 4.26 -0.511
DAVOL10 -7.44 -6.04 54.1 -0.504 -0.071 0.920 45.45 0.248 3.38 -0.812
DAVOL20 -10.95 -9.78 54.4 -0.702 -0.119 0.919 48.96 0.246 2.86 -1.079
DAVOL5 -9.54 -8.28 51.8 -0.612 -0.097 0.950 46.41 0.250 3.17 -1.019
EMA5 2.77 4.85 48.3 -0.047 0.049 1.041 30.70 0.259 4.66 -0.068
EMAC10 0.27 2.17 50.6 -0.144 0.024 1.037 31.42 0.258 4.19 -0.203
EMAC12 1.58 3.57 50.6 -0.093 0.037 1.032 31.76 0.258 4.17 -0.131
EMAC20 6.01 8.29 54.1 0.069 0.079 1.044 32.43 0.260 4.32 0.102
EMAC26 3.83 5.97 54.1 -0.008 0.059 1.046 34.87 0.261 4.29 -0.012
EMAC120 1.74 3.75 48.3 -0.088 0.034 0.960 35.81 0.254 3.77 -0.145
Kurtosis20 -2.96 -1.27 52.5 -0.395 -0.027 0.772 27.71 0.182 2.74 -0.6
Kurtosis60 -0.28 1.59 53.9 -0.240 -0.001 0.716 22.07 0.179 2.64 -0.367
Kurtosis120 -4.12 -2.51 49.0 -0.468 -0.043 0.708 28.28 0.182 2.44 -0.684
MAC5 0.54 2.46 46.9 -0.134 0.026 1.035 30.03 0.257 4.29 -0.195
MAC10 -1.20 0.61 49.6 -0.206 0.008 1.039 34.15 0.256 4.19 -0.286
MAC20 2.47 4.52 53.5 -0.059 0.046 1.043 34.58 0.257 4.09 -0.083
MAC60 11.10 13.72 53.8 0.243 0.120 0.987 28.61 0.254 4.34 0.399
MAC120 3.11 5.21 47.7 -0.035 0.049 0.987 35.40 0.259 3.99 -0.058
MACDC -11.12 -9.97 52.4 -0.674 -0.117 1.003 46.48 0.261 3.54 -1.046
MASS -10.55 -9.35 51.0 -0.673 -0.115 0.886 48.34 0.249 6.19 -1.032
MAWVAD 7.29 9.67 55.9 0.119 0.082 0.891 36.54 0.248 4.19 0.184
MFI14 -6.21 -4.73 51.5 -0.456 -0.056 0.918 45.65 0.241 3.61 -0.699
PLRC6 -14.83 -13.92 47.8 -0.905 -0.178 0.960 66.24 0.259 3.85 -1.368
PLRC12 -7.28 -5.87 53.3 -0.452 -0.064 1.012 52.30 0.273 4.20 -0.682
PLRC24 -5.77 -4.26 58.1 -0.361 -0.044 1.034 53.28 0.289 4.88 -0.549
PSY -9.11 -7.83 50.6 -0.660 -0.096 0.880 49.79 0.223 3.36 -0.992
Price1M -13.55 -12.55 53.0 -0.760 -0.156 0.977 64.85 0.280 4.21 -1.145
Price3M -6.69 -5.25 58.4 -0.380 -0.056 1.023 59.14 0.305 5.44 -0.555
Price1Y 1.55 3.54 66.8 -0.075 0.041 1.116 60.25 0.326 7.28 -0.105
ROC6 -16.13 -15.30 47.4 -0.982 -0.200 0.978 68.14 0.262 3.56 -1.741
ROC20 -11.63 -10.51 54.7 -0.628 -0.123 1.014 60.28 0.291 4.51 -0.971
ROC60 -6.86 -5.43 59.6 -0.382 -0.056 1.062 62.16 0.309 5.28 -0.529
ROC120 2.36 4.41 66.9 -0.051 0.049 1.117 56.36 0.319 6.79 -0.071
Skewness20 -5.91 -4.41 50.8 -0.538 -0.058 0.820 40.65 0.197 3.04 0.839
Skewness60 -6.47 -5.01 51.9 -0.595 -0.068 0.765 38.02 0.190 2.58 -0.917
Skewness120 -4.37 -2.77 51.1 -0.459 0.044 0.743 36.01 0.192 2.69 -0.684
TRIX5 -10.72 -9.54 52.5 -0.607 -0.110 1.009 56.07 0.280 4.16 -0.923
TRIX10 -3.51 -1.85 58.4 -0.268 -0.019 1.008 52.92 0.292 5.14 -0.4
TVMA20 -1.92 -0.16 64.9 -0.209 0.014 1.268 44.66 0.288 4.63 -0.31
TVMA6 -0.75 1.09 62.4 -0.166 0.026 1.259 39.46 0.288 4.76 -0.246
TVSTD20 -1.83 -0.06 58.5 -0.214 0.012 1.218 35.89 0.277 4.03 -0.318
TVSTD6 -3.77 -2.13 53.4 -0.297 -0.010 1.202 46.57 0.272 4.40 -0.449
VDEA -3.05 -1.37 57.6 -0.305 -0.023 0.857 37.01 0.239 3.06 -0.484
VDIFF -7.93 -6.57 0.2 -0.533 -0.079 0.895 39.79 0.247 3.08 -0.842
VEMA5 -0.41 1.46 57.9 -0.186 0.006 0.850 37.31 0.237 3.33 -0.277
VEMA10 0.95 2.91 58.9 -0.128 0.019 0.844 35.51 0.236 3.36 -0.19
VEMA12 1.01 2.96 59.3 -0.126 0.020 0.840 35.17 0.235 3.35 -0.186
VEMA26 2.41 4.46 59.3 -0.068 0.033 0.838 31.03 0.231 3.37 -0.1
VMACD -1.03 0.79 50.5 -0.216 0.001 0.875 33.96 0.235 3.38 -0.345
VOL5 -8.12 -6.76 55.9 -0.452 -0.066 1.155 53.86 0.296 4.48 -0.656
VOL10 -7.07 -5.64 58.3 -0.404 -0.053 1.145 50.65 0.298 4.35 -0.59
VOL20 -5.43 -3.89 58.6 -0.337 -0.032 1.156 50.19 0.298 4.33 -0.481
VOL60 0.42 2.34 58.1 -0.125 0.032 1.145 38.35 0.286 4.71 -0.175
VOL120 -1.88 -0.12 53.0 -0.213 0.007 1.136 42.09 0.281 4.64 -0.3
VOL240 -5.51 -3.98 47.0 -0.382 -0.038 1.075 43.53 0.265 3.99 -0.541
VOSC -6.35 -4.88 51.2 -0.453 -0.056 0.944 40.40 0.246 3.44 -0.739
VROC6 -16.30 -15.49 42.7 -1.179 -0.208 0.897 69.03 0.221 2.91 -1.778
VROC12 -11.47 -10.34 46.7 -0.793 -0.127 0.915 51.62 0.228 2.74 -1.241
VR -8.43 -7.10 52.8 -0.566 -0.083 0.932 48.57 0.244 3.72 -0.834
VSTD10 -0.77 1.07 56.3 -0.205 0.001 0.843 37.75 0.234 3.19 -0.309
VSTD20 -1.60 0.19 53.7 -0.245 -0.007 0.844 37.31 0.232 3.01 -0.366
Variance20 -1.94 -0.18 59.9 -0.201 0.005 1.124 46.38 0.302 4.84 -0.301
Variance60 9.96 12.51 65.2 0.168 0.122 1.178 30.63 0.313 5.91 0.241
Variance120 5.40 7.65 60.8 0.040 0.083 1.193 39.54 0.312 5.93 0.058
Volume1M -15.54 -14.68 20.7 -0.867 -0.189 0.991 63.68 0.285 3.61 -1.339
WVAD -0.71 1.13 52.3 -0.199 0.002 0.836 42.38 0.238 3.67 -0.317
arron_down_25 -7.35 -5.94 48.3 -0.593 -0.072 0.880 42.63 0.209 2.84 -0.826
arron_up_25 -15.51 -14.65 46.4 -1.049 -0.195 0.870 66.73 0.235 3.19 -1.604
bear_power -13.31 -12.30 50.8 -0.800 -0.154 0.934 59.58 0.261 3.61 -1.182
beta 10.30 12.87 62.4 0.168 0.134 1.344 35.96 0.328 6.76 0.252
book_to_price 1.87 3.86 58.5 -0.117 0.031 0.889 26.73 0.181 2.91 -0.162
bull_power -14.81 -13.91 51.1 -0.844 -0.176 0.984 65.99 0.278 3.83 -1.285
earnings_yield -2.41 -0.68 52.6 -0.348 -0.029 0.627 25.87 0.188 2.55 -0.55
growth 2.90 4.99 59.9 -0.045 0.052 1.068 37.35 0.244 4.50 -0.067
leverage 2.05 4.07 53.1 -0.089 0.023 0.719 22.57 0.217 3.22 -0.144
liquidity -2.71 -1.00 56.7 -0.243 -0.002 1.137 41.28 0.284 4.56 -0.35
momentum 10.93 13.54 65.8 0.195 0.129 1.159 43.19 0.310 6.82 0.271
money_flow_20 -1.72 0.05 65.5 -0.202 0.016 1.269 44.10 0.288 4.63 -0.298
price_no_fq 6.83 9.17 76.1 0.095 0.092 1.132 44.86 0.267 6.20 0.131
pull-up 2.19 4.23 54.5 -0.068 0.044 1.062 33.09 0.265 3.92 -0.093
pull-down -2.28 -0.54 47.6 -0.317 -0.015 0.831 34.46 0.203 3.87 -0.439
residual_volatility 4.13 6.30 62.3 0.003 0.062 1.048 41.34 0.269 5.25 0.004
sharpe_ratio_20 -11.88 -10.78 54.0 -0.663 -0.129 0.979 57.66 0.282 4.09 -1.019
sharpe_ratio_60 -7.42 -6.02 58.8 -0.417 -0.065 1.029 65.19 0.300 5.17 -0.583
sharpe_ratio_120 -0.32 1.55 65.7 -0.143 0.019 1.062 60.52 0.303 6.10 -0.2
size -6.36 -4.89 51.8 -0.571 -0.058 0.906 36.00 0.196 2.64 -0.82
turnover_volatility -10.80 -9.62 51.8 -0.624 -0.108 1.075 57.05 0.274 3.70 -0.905
1day_VPT -16.45 -15.65 44.8 -1.062 -0.211 0.904 70.43 0.248 3.09 -1.553
1day_VPT_6 -6.36 -4.89 53.0 -0.416 -0.053 0.995 47.06 0.269 3.81 -0.82
1day_VPT_12 -0.07 1.82 57.9 -0.146 0.020 1.033 46.51 0.278 4.39 -0.22
25week_close -11.51 -10.39 51.0 -0.859 -0.133 0.817 47.79 0.211 2.44 -0.684
QF_Month_1 17.35 19.43 57.8 0.915 0.162 0.588 18.35 0.161 2.81 0.946
QF_Month_2 9.91 13.53 49.3 0.289 0.102 0.684 26.77 0.205 3.61 0.372
QF_Month_3 7.37 10.4 51.6 0.246 0.064 0.483 17.33 0.137 2.3 0.309
QF_Week_1 -0.83 1.41 46.4 -0.236 -0.003 0.726 29.31 0.204 3.05 -0.291
QF_Week_2 7.49 10.55 49.2 0.16 0.085 0.792 29.08 0.219 3.73 0.206
QF_Week_3 12.3 12.4 54.4 0.372 0.116 0.793 29.86 0.223 3.77 0.48
QF_Day_1 7.89 11.4 43.1 0.181 0.09 0.834 29.15 0.196 3.92 0.245
QF_Day_2 10.23 10.68 44.4 0.279 0.097 0.819 30.02 0.223 3.91 0.361
QF_Day_3 9.81 10.03 44.6 0.284 0.098 0.801 30.1 0.223 1.02 0.367
QF_1013.12 17.73 61.8 0.754 0.128 0.594 18.32 0.159 2.02 0.714
QF_512.59 16.02 61 0.63 0.117 0.504 13.3 0.136 2.01 0.796
QF_124.71 35.34 53.3 0.967 0.249 0.667 23.28 0.214 3.35 1.299
Average -2.92 -1.25 53.28 -0.32 -0.02 0.96 43.61 0.25 3.99 -0.467
CSI300 1.77 \ \ 0.009 \ \ \ 0.197 3.19 \
Appendix B Comparison between different quantformer strategies
(a) Factor performance
(b) Comparison between models
Figure 6: Comparison of factor performance and model performance
In the provided scatter plot 6(a), the $QF\_Month\_1$ factor exhibits notable performance in terms of annual return. Compared to other factors, $QF\_Month\_1$ offers better returns. Regarding risk, as measured by the $Variance\_Covariance\_VaR\_99$, the $QF\_Month\_1$ factor’s data points are positioned in a moderately high range, indicating a balanced risk profile. The size and color intensity of the $QF\_Month\_1$ points, reflecting higher Alpha values, suggest a robust excess return or relative performance. Combining these observations, the $QF\_Month\_1$ factor likely provides a compelling investment opportunity by delivering robust returns while maintaining a moderate risk level, an attractive proposition for strategies aiming to optimize the trade-off between risk and return.
In scatter plot 6(b), most of the factors based on quantformer perform better compared with the other 100 factors. The performance of quantformer factors which are trained under different scales shows ability in making profit and risk management.